6.02×10 20 molecules of urea are present in 100ml of its solution the cocentration of the solution is

6.02×1020molecules of urea are present in 100ml of its solution the cocentration of the solution is 
 

Grade:12th pass

2 Answers

varun
6 years ago
HELLO,
Moles of urea = molecules/avogadro no
                       =    6.02X10^20 / 6.022 X 10^23
                       =          10^-3 moles
So it will have 10^-2 mole in 1 litre of solution
because 
Molarity/concentration = mole / litre
                                         = 10^-3/1L
                                        =    10^-2M
THANK YOU
Yash Chourasiya
3 years ago
Dear Student

Molarity = molesofsolute​ /volumeofsolutioninlitre
NA​ = 6.02×1023
Number of moles = ​Numberofparticle/NA
= 6.02×1020​/6.023×1023
Number of moles=0.001

Molarity = 0.001/0.1​=0.01M

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya